A particular manufacturing design requires a shaft with a diameter between 19.89 mm and 20.013 mm. The manufacturing process yields shafts with diameters normally distributed, with a mean of 20.002 mm and a standard deviation of 0.005 mm. Complete parts (a) through (c). a. For this process what is the proportion of shafts with a diameter between 19.89 mm and 20.00 mm? The proportion of shafts with diameter between 19.89 mm and 20.00 mm is Round to four decimal places as needed.)
Answer #1

Solution :

Given that ,

mean = \mu = 20.002

standard deviation = \sigma = 0.005

P(19.89 < x < 20.0 )

= P((19.89 - 20.002)/ 0.005) < (x - \mu ) / \sigma < (20.00 - 20.002 ) / 0.005) )

= P(-22.4 < z < -0.4)

= P(z < -0.4) - P(z < -22.4)

= 0.3446 - 0

= 0.3446

Proportion = 0.3446

Probability =